OpenStudy (anonymous):

Solve the system using substitution. What is the x- or y-coordinate of the solution? 2x - 3y = -1 x - 2y = -5

OpenStudy (anonymous):

these are equations of lines and you will get a solution if they intersect. You can try to plot by hand and see the point of intersection. For substitution, you can see that the second equation can give you x by simply adding 2y on both sides. Use that value in the first equation. Will that work for you?

OpenStudy (anonymous):

Which One Would Be The Answer ? A.-11 B.-10 C.9 D.11

OpenStudy (anonymous):

@rkopen

OpenStudy (anonymous):

x = 2y+5 use this in (1) so, 2(2y+5) -3y = -1 4y + 10 -3y = -1 y + 10 = -1 y = 11 now plug this into x = 2y + 5 x = 22 + 5 = 27

OpenStudy (anonymous):

Oh, my x value was wrong x = 2y - 5 can you work from there?